Solve the following equations.   2x/3 = 18 
Solve the following equations.   2x/3 = 18 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

2x/3 = 18
2x=18×3
2x=54
x=54/2
x=27

Thanks